OpenStudy (anonymous):

Jack spent ¹3/4 hors doing homework. 1/2 hours eating. And ¹1/4 playing video games. Jack got home from school at 4:30pm. what time did he finish all three tasks?

OpenStudy (aum):

Add \(\large \frac 34 + \frac 12 + \frac 14 \). The first and the last fractions are easier to add. Then add the middle fraction. Add the result to 4:30 PM to find the time when all three tasks are done.